# Hypergraph clustering using Ricci curvature: an edge transport perspective

### v1

@@ Abstract
Keywords: Hypergraph, Clustering, Ricci Curvature, Optimal Transport
@@ 1 Introduction
Hypergraphs are a generalization of graphs extending the concept of edges to encompass relationships involving any number of vertices rather than being restricted to pairs. Hypergraphs provide a more faithful data representation than graphs as many real-life interactions are at a group level instead that at a binary one. Examples include social interactions, simultaneous participation at an event, co-authorship, mutual evolutionary traits, spatial proximity or molecular interactions. These interactions are better encoded by hypergraphs instead of traditional graphs, see Davis (1941); Dotson et al. (2022); Konstantinova and Skorobogatov (2001); Torres et al. (2021); Klamt et al. (2009). We refer to Berge (1984) and Bretto (2013) for a comprehensive overview of hypergraph theory.
@@
Recently, machine learning tasks on hypergraphs have encountered a growing success, in particular for classification (Feng et al., 2019), regression (Liu et al., 2021), anomaly detection (Lee et al., 2022) or embedding (Zhang et al., 2019; Antelmi et al., 2023). This article focuses on the clustering problem or community detection, i.e. recovering labels on the nodes in a fully unsupervised setting from a single hypergraph instance. Several approaches have been taken to tackle this problem such as embeddings using neural networks (Zhang et al., 2021; Lee and Shin, 2023), random walks-based embeddings (Hayashi et al., 2020; Huang et al., 2019) followed by a clustering in the Euclidean space. Some methods such as (Kumar et al., 2020; Kamiński et al., 2024) aim at maximizing hypergraph-modularity functions that measure the strength of a clustering. Finally, let us mention the work of Eyubov et al. (2023) that finds a partition in linear time by streaming all the nodes one by one.
@@
The present work investigates generalizations of Ricci flow based clustering on hypergraphs. Ollivier-Ricci curvature, introduced by Ollivier (2007) for metric spaces and later adapted to graphs (Lin et al., 2011), defines a distance on graph edges that quantifies local curvature using optimal transport theory. Edges within strongly connected communities have a positive curvature while edges bridging two communities are negatively curved. Ricci curvature can be turned into a flow dynamic in order to further stress out this community structure and take some more global graph properties into account. Ricci curvature on graphs has been used to derive theoretical bounds on the spectrum of the graph Laplacian by Lin and Yau (2010) and Bauer et al. (2011). For applied purposes, it has been used for data exploration (Ni et al., 2015), representation learning (Zhang et al., 2023), topological data analysis (Carriere and Blumberg, 2020; Hacquard and Lebovici, 2024), and in particular clustering (Ni et al., 2019), relying on the property of the Ricci flow to place emphasis on the community structure.
@@
There has been a few attempts to generalize Ollivier-Ricci curvature to hypergraphs, in particular by Coupette et al. (2022). The authors introduce a notion of curvature using random walks on the nodes. As pointed out by Chitra and Raphael (2019), in most practical cases, considering a random walk on the hypergraph is equivalent to a random walk on its clique expansion (a graph representation where each hyperedge is replaced by a weighted clique). Doing so is a common way to circumvent the complex hypergraph structure by replacing it with a graph with a similar structure. However, as noted in Chitra and Raphael (2019), this reduction can result in significant information loss, as distinct hypergraphs may share identical clique graphs (Hein et al., 2013). In addition, replacing a single hyperedge between $k$ nodes implies adding up to (${k\choose 2}$ connections between nodes which can be computationally prohibitive in a network with mutual interactions between many agents. Directed hypergraph extensions of Ricci curvature have also been proposed (Eidi and Jost, 2020).
@@
The main contribution of this works aims at defining a notion of Ricci curvature on hypergraphs where we consider probability distributions on the edges instead than on the nodes. This approach leverages the line expansion of the hypergraph, which is a graph representation where nodes correspond to hyperedges, and edges reflect hyperedges intersections. The line expansion is a common way to represent a hypergraph for learning purposes, see Bandyopadhyay et al. (2020); Yang et al. (2022). In addition, it has been demonstrated in Kirkland (2018) that almost all the hypergraph information is retained by the joint knowledge of both its clique and line expansions. The current work puts a particular focus on clustering. We perform a thorough experimental study on synthetic and real data comparing the approach where we transport measures on the edges to the more standard approach where we transport measures on the nodes. More precisely, this notion of edge transport should be favored when we have a large number of small communities, when the edges between communities have a larger cardinality than the ones within communities, and more generally for hypergraphs with very large hyperedges as this approach is much more efficient computationally.
@@ 2.1 Hypergraphs and associated graphs
Undirected hypergraphs, sometimes referred to as hypernetworks or 2-modes networks consist of a set of nodes $V$ and a set of hyperedges $E$. Extending the concept of graphs where edges link two distinct nodes, hyperedges are defined as non-empty sets of nodes of arbitrary size. For a hyperedge $e=(u_{1},\ldots,u_{k})$, its cardinality $k$ is referred to as its size. A hypergraph in which all hyperedges have the same size $k$ is called a $k-$uniform hypergraph. Notably, graphs are $2-$uniform hypergraphs. We also give ourselves a weight function $w:E\to\mathbb{R}_{\geq 0}$. For the sake of simplicity and to only focus on the topological structure of the hypergraphs, all further definitions assume unweighted hypergraphs, i.e. $w$ constant equal to 1. We further define the star of a vertex $v\in V$ as the set of hyperedges that include $v$: $St(v)=\{e\in E|v\in e\}$.
@@
Given a hypergraph $H$ with node set $V=(v_{1},\ldots,v_{n})$ and edge set $E=(e_{1},\ldots,e_{m})$, its incidence matrix $\mathbf{I}\in\{0,1\}^{n\times m}$ is defined as $\mathbf{I}_{i,j}=1$ if $v_{i}\in e_{j}$ and 0 otherwise. This representation as a rectangular matrix makes further analysis of hypergraphs much more complicated than that of graphs, which can be described by their adjacency matrix. A commonly used simplification of hypergraphs is the clique expansion, where each edge is replaced by a clique. More precisely, the clique expansion $\mathbf{C}(H)$ of the hypergraph $H$ is the graph with node set $V$ and with an edge $(x,y)$ if there exists a hyperedge $e\in E$ such that $x,y\in e$. The clique expansion provides a convenient simplification of hypergraphs by grouping nodes belonging to the same edge. However, there is a certain loss of information since multiple hypergraphs share the same clique expansion, see Figure 1. We refer to Hein et al. (2013) for a more thorough analysis of differences between hypergraphs and their clique expansion. Notably, according to Chitra and Raphael (2019) and Hayashi et al. (2020), unless the nodes are assined an edge-dependent weighting, standard notions of random walks and Laplacians on the hypergraphs can be expressed on a reweighted clique expansions. This implies that directly applying Laplacian-based methods, such as clustering, on hypergraphs often incurs the information loss inherent in their clique expansions. We note that the adjacency matrix $A_{C}$ of the clique graph can be derived from the incidence matrix via $A_{C}=\mathbf{I}\mathbf{I}^{T}-D_{V}$ where $D_{v}$ is the diagonal node-degree matrix, with entries $d_{i}=\sum_{j=1}^{n}\mathbf{I}_{i,j}$.
@@
Another graph associated with $H$ is the line graph $\mathbf{L}(H)$. Each edge in $H$ now corresponds to a node in $\mathbf{L}(H)$. Thus, $\mathbf{L}(H)$ has node set $(e_{1}^{\star},\ldots,e_{m}^{\star})$ and we have an edge $(e_{i}^{\star},e_{j}^{\star})$ in $\mathbf{L}(H)$ if the corresponding hyperedges intersect in $H$, i.e. $e_{i}\cap e_{j}\neq\emptyset$. Similarly to the clique expansion, many hypergraphs share the same line expansion, see Figure 1. However, joint knowledge of both the line and clique expansions seems to capture most of the structural information of the hypergraph. Indeed in Figure 1, hypergraphs $H_{1}$, $H_{2}$ and $H_{3}$ all share the same clique expansion while having different line expansions (they even have different node sets). Similarly to the clique graph, the adjacency matrix $A_{L}$ of the line graph can be derived from the incidence matrix via $A_{L}=\mathbf{I}^{T}\mathbf{I}-D_{e}$ where $D_{e}$ is the diagonal matrix of edge-lengths, with entries $\delta_{i}=|e_{i}|$. Let us define the dual hypergraph $H^{\star}=(E^{\star},V^{\star})$ as the hypergraph obtained by swapping the edge and node sets, i.e. $E^{\star}=(e_{1}^{\star},\ldots,e_{m}^{\star})$ and $V^{\star}=(v_{1}^{\star},\ldots v_{n}^{\star})$ where $v_{i}^{\star}=\{e_{j}^{\star}|v_{i}\in e_{j}\}$. We can easily observe, see Zhou et al. (2022), that the clique expansion is the line graph of the dual hypergraph: $\mathbf{C}(H)=\mathbf{L}(H^{\star})$.
@@
We can wonder if the joint knowledge of both the clique graph and the line graph allows to reconstruct the hypergraph. The answer is negative, as there exists pairs of distinct 0-1 matrices $(A,B)$ such that $AA^{T}=BB^{T}$ and $A^{T}A=B^{T}B$. Such matrices are called Gram mates, and we refer to Kim and Kirkland (2022) for proof of this result and a thorough analysis of families of matrices that yield Gram mates. However, it has been demonstrated in Corollary 1.1.1 of Kirkland (2018) that if we uniformly pick two $m\times n$ 0-1 matrices, the probability that they are Gram mates decays exponentially as $(m,n)\to\infty$. Thus, for practical applications, the information loss from replacing a hypergraph with its clique and line expansions together is negligible. This representation also simplifies hypergraph analysis significantly.
@@
We consider hypergraphs where nodes are associated with discrete labels $\{1,\ldots,K\}$, representing communities or clusters. Intuitively, such hypergraphs exhibit a community structure, with denser connections between nodes sharing the same label compared to those with different labels. We assume a fully unsupervised setting, where we try to infer the labels (up to permutation) simply via the hypergraph structure. This generalization of the graph partitioning problem has been studied extensively, and we refer to Çatalyürek et al. (2023) for a survey of different approaches to tackle this problem. A popular method to determine the quality of a partitioning is given by modularity maximization, initially defined by Newman (2006). We recall its definition in the graph case. For a graph $G=(V,E)$ and a partitioning $\mathbf{C}=(C_{1},\ldots,C_{K})$ of $V$ into $K$ subsets, we define the modularity function $Q$ as:
$Q(\mathbf{C})=\sum_{i=1}^{K}\frac{e_{G}(C_{i})}{|E|}-\sum_{i=1}^{K}\left(\frac{vol(C_{i})}{vol(V)}\right)^{2},$
where $e_{G}(C_{i})$ is the number of edges in the subgraph of $G$ induced by a node set $C_{i}$, and $vol(A)=\sum_{v\in A}deg(v)$ is the volume of a subset $A$ of nodes. Up to renormalization, the modularity function computes the difference between the number of edges uncut by the partitioning and the expected number of edges the same partitioning would yield in the Chung-Lu random graph model introduced in Chung and Lu (2006). Maximizing $Q$ over every partitioning yields partitions that minimize cuts within communities, aligning with the underlying structure. There have been several works generalizing modularity for hypergraphs, notably Kumar et al. (2020) and Kamiński et al. (2024), by generalizing the Chung-Lu random model to hypergraphs, and allowing different weights to hyperedges of different sizes. Modularity also serves as a ground truth-free measure of the quality of a given clustering.
@@ 2.2 Ricci curvature and flows on graphs
This paper addresses the clustering problem using an extension of Ollivier-Ricci curvature to hypergraphs. This method extends prior works on Ollivier-Ricci curvature, initially defined in Ollivier (2007) on metric spaces, and extended to the case of graphs in Lin et al. (2011). In Ni et al. (2019), the Ricci curvature is turned into a discrete flow such that edges between communities are heavily curved. They directly derive a clustering algorithm. We start by recalling the key concepts of their method in the case of graphs.
@@
Let $G$ be a weighted graph with node set $V=(x_{1},\ldots,x_{n})$, edge set $E$ and a weight function $w:E\to\mathbb{R}_{\geq 0}$. In what follows, the function $w$ is set to be a dissimilarity measure between nodes. Let $d$ denote the shortest path distance between nodes induced by weights $w$. For each node $x$, we define a probability measure $\nu_{x}$ over $x$ and its neighbors given by:
$\nu_{x}(y)=\begin{cases}\alpha&\text{if }y=x,\\ \frac{1-\alpha}{|\mathcal{N}(x)|}&\text{if }y\in\mathcal{N}(x),\\ 0&\text{ otherwise, }\end{cases}$ (1)
where $\mathcal{N}(x)$ denotes the set of neighbors of $x$ and $\alpha\in[0,1]$ is a real parameter controlling the amount of mass retained at $x$. More general measures on the nodes are considered in Ni et al. (2019).
@@
where $W$ is the Wasserstein distance between probability measures for the cost function $d$, see Santambrogio (2015). More precisely, a discrete transport plan $T$ between the measures $\nu_{x}$ and $\nu_{y}$ is an application $T:V\times V\to[0,1]$ such that $\sum_{v^{\prime}\in V}A(u,v^{\prime})=\nu_{x}(u)$ and $\sum_{u^{\prime}\in V}A(u^{\prime},v)=\nu_{y}(v).$ The Wasserstein distance corresponds to the total cost of moving $\nu_{x}$ to $\nu_{y}$ with the optimal transport plan for a cost function $d$, i.e.
$W(\nu_{x},\nu_{y})=\inf\left\{\sum_{u,v\in V}T(u,v)d(u,v)|T\text{ transport plan between }\nu_{x}\text{ and }nu_{y}\right\}.$
The sign of $\kappa$ provides information on the density of connections. Indeed, two nodes $x$ and $y$ belonging to the same community will typically share many neighbors, hence transporting $\nu_{x}$ to $\nu_{y}$ has a smaller cost than moving $x$ to $y$ and therefore induces a positive curvature. Conversely, nodes belonging to different communities do not share many neighbors, such that moving $\nu_{x}$ to $\nu_{y}$ requires traveling through the edge $(x,y)$ bridging the two communities. This in turn corresponds to a negative curvature. We refer to Figure 2 for a concrete example: when transporting $\nu_{v_{i}}$ to $\nu_{v_{j}}$ and $\nu_{u_{i}}$ to $\nu_{u_{j}}$, most of the mass remains on the nodes such that measures are transported at a low cost. If $\alpha=1/5$, all the mass remains and $\kappa(v_{i},v_{j})=\kappa(u_{i},u_{j})=1$. When transporting $\nu_{v_{1}}$ to $\nu_{u_{1}}$, all the $v_{i}$’s must be transported to $u_{i}$’s through the edge $(u_{1},v_{1})$ such that the total transport cost is larger than $d(u_{1},v_{1})$.
@@
It instantly appears that trimming edges of low curvature naturally partitions the graph where each connected component corresponds to a different community. We note that Ricci curvature can theoretically be computed for any pair of nodes in the graph. However, for our purposes, we only need to compute it for adjacent nodes.
@@
where $\kappa^{(l)}$ is the Ricci curvature computed with the distance induced by the weights $w^{(l)}$. The updated weights after the $l$-th iteration are referred to as the Ricci flow. This flow dynamic stretches edges of low curvature, making it more and more costly to transport measures between communities. The complete clustering algorithm is described by Algorithm 1, and simply trims edges with a high Ricci flow. The choice of the number of flow iterations $N$ is mostly heuristic, we refer to (Ni et al., 2019) and the corresponding python library GraphRicciCurvature for some guidelines. The threshold $\tau$ at which to trim the weights is a crucial parameter and can also be taken as the one that maximizes the modularity function defined in Section 2.1, which evaluates clustering quality without ground truth labels.
Input: Graph $G=(V,E)$, Number of iterations $N$, Threshold $\tau$
@@
This approach ensures that inter-community connections are down-weighted while preserving intra-community coherence, enabling a robust and interpretable clustering.

### v2

@@ 1 Introduction
Hypergraphs are generalizations of graphs extending the concept of edges to encompass relationships involving any number of vertices rather than being restricted to pairs. Hypergraphs provide a more faithful data representation than graphs as many real-life interactions occur at a group level rather than at a binary one. Examples include social interactions, simultaneous participation at an event, co-authorship, mutual evolutionary traits, spatial proximity or molecular interactions. These interactions are better encoded by hypergraphs instead of traditional graphs, see Davis, (1941); Dotson et al., (2022); Konstantinova and Skorobogatov, (2001); Torres et al., (2021); Klamt et al., (2009). We refer to Berge, (1984) and Bretto, (2013) for a comprehensive overview of hypergraph theory.
@@
Recently, machine learning tasks on hypergraphs have encountered a growing success, in particular for classification (Feng et al.,, 2019), regression (Liu et al.,, 2021), anomaly detection (Lee et al.,, 2022) or embedding (Zhang et al.,, 2019; Antelmi et al.,, 2023). The present article focuses on the clustering problem or community detection, i.e. recovering labels on the nodes in a fully unsupervised setting from a single hypergraph instance. Several approaches have been taken to tackle this problem such as embeddings using neural networks (Zhang et al.,, 2021; Lee and Shin,, 2023), random walks-based embeddings (Hayashi et al.,, 2020; Huang et al.,, 2019) followed by a clustering in the Euclidean space. Some methods developed in Kumar et al., (2020); Kamiński et al., (2024) aim at maximizing hypergraph-modularity functions that measure the strength of a clustering. Finally, let us mention the work of Eyubov et al., (2023) that finds a partition in linear time by streaming all the nodes one by one.
@@
The present work investigates generalizations of Ricci flow based clustering on hypergraphs. Ollivier-Ricci curvature, introduced by Ollivier, (2007) for general metric spaces and later specifically tailored to graphs (Lin et al.,, 2011), defines a distance on graph edges that quantifies local curvature using optimal transport theory. Edges within strongly connected communities have a positive curvature while edges bridging two communities are negatively curved. Ricci curvature can be turned into a flow dynamic in order to further stress out this community structure and take some more global graph properties into account. Alternative notions of curvature for graph processing are also present in Forman, (2003); Sreejith et al., (2016); Iyer et al., (2024); Tian et al., (2025). Ricci curvature on graphs has been used to derive theoretical bounds on the spectrum of the graph Laplacian by Lin and Yau, (2010) and Bauer et al., (2011). For applied purposes, it has been used for data exploration (Ni et al.,, 2015), representation learning (Zhang et al.,, 2023), topological data analysis (Carriere and Blumberg,, 2020; Hacquard and Lebovici,, 2024), and in particular clustering (Ni et al.,, 2019), relying on the property of the Ricci flow to place emphasis on the community structure.
@@
There has been a few attempts to generalize Ollivier-Ricci curvature to hypergraphs, in particular by Coupette et al., (2022). The authors introduce a notion of curvature using random walks on the nodes. As pointed out by Chitra and Raphael, (2019), in most practical cases, considering a random walk on the hypergraph is equivalent to a random walk on its clique expansion (a graph representation where each hyperedge is replaced by a weighted clique). Doing so is a common way to circumvent the complex hypergraph structure by replacing it with a graph with a similar structure. However, as noted in Chitra and Raphael, (2019), this reduction can result in significant information loss, as distinct hypergraphs may share identical clique graphs (Hein et al.,, 2013). In addition, replacing a single hyperedge between $k$ nodes implies adding up to $k\choose 2$} connections between nodes which can be computationally prohibitive in a network with mutual interactions between many agents. Extensions of Ricci curvature on directed hypergraphs have also been proposed by Eidi and Jost, (2020).
@@
The main contribution of this work aims at defining a notion of Ricci curvature on hypergraphs where we consider probability distributions on the edges instead than on the nodes. This approach leverages the line expansion of the hypergraph, which is a graph representation where nodes correspond to hyperedges, and edges reflect hyperedges intersections. The line expansion is a common way to represent a hypergraph for learning purposes, see Bandyopadhyay et al., (2020); Yang et al., (2022). In addition, it has been demonstrated in Kirkland, (2018) that almost all the hypergraph information is retained by the joint knowledge of both its clique and line expansions. The current work puts a particular focus on clustering. We perform a thorough experimental study on synthetic and real data comparing the approach where we transport measures on the edges to the more standard approach where we transport measures on the nodes. More precisely, this notion of edge transport should be favored when we have a large number of small communities, when the edges between communities have a larger cardinality than the ones within communities, and more generally for hypergraphs with very large hyperedges as this approach is much more efficient computationally.
@@ 2.1 Hypergraphs and associated graphs
Undirected hypergraphs, sometimes referred to as hypernetworks or 2-modes networks, consist of a set of nodes $V$ and a set of hyperedges $E$. Extending the concept of graphs where edges link two distinct nodes, hyperedges are defined as non-empty sets of nodes of arbitrary size. For a hyperedge $e=(u_{1},\ldots,u_{k})$, its cardinality $k$ is referred to as its size. A hypergraph in which all hyperedges have the same size $k$ is called a $k$-uniform hypergraph. Notably, graphs are 22-uniform hypergraphs. We further define the star of a vertex $v\in V$ as the set of hyperedges that include $v$: $St(v)=\{e\in E|v\in e\}$.
@@
Given a hypergraph $H$ with node set $V=(v_{1},\ldots,v_{n})$ and edge set $E=(e_{1},\ldots,e_{m})$, its incidence matrix $\mathbf{I}\in\{0,1\}^{n\times m}$ is defined as $\mathbf{I}_{i,j}=1$ if $v_{i}\in e_{j}$ and 0 otherwise. This representation as a rectangular matrix makes further analysis of hypergraphs much more complicated than that of graphs, which can be described by their adjacency matrix. A commonly used simplification of hypergraphs is the clique expansion, where each edge is replaced by a clique. More precisely, the clique expansion $\mathbf{C}(H)$ of the hypergraph $H$ is the graph with node set $V$ and with an edge $(x,y)$ if there exists a hyperedge $e\in E$ such that $x,y\in e$. The clique expansion provides a convenient simplification of hypergraphs by grouping nodes belonging to the same edge. However, there is a certain loss of information since multiple hypergraphs share the same clique expansion, see Figure 1. We refer to Hein et al., (2013) for a more thorough analysis of differences between hypergraphs and their clique expansion. Notably, according to Chitra and Raphael, (2019) and Hayashi et al., (2020), unless the nodes are assigned an edge-dependent weighting, standard notions of random walks and Laplacians on the hypergraphs can be expressed on a reweighted clique expansion. This implies that directly applying Laplacian-based methods, such as clustering, on hypergraphs often incurs the information loss inherent in their clique expansions. We note that the adjacency matrix $A_{C}$ of the clique graph can be derived from the incidence matrix via $A_{C}=\mathbf{I}\mathbf{I}^{T}-D_{V}$ where $D_{v}$ is the diagonal node-degree matrix, with entries $d_{i}=\sum_{j=1}^{n}\mathbf{I}_{i,j}$.
@@
Another graph associated with $H$ is the line graph $\mathbf{L}(H)$. Each edge in $H$ now corresponds to a node in $\mathbf{L}(H)$. Thus, $\mathbf{L}(H)$ has node set $(e_{1}^{\star},\ldots,e_{m}^{\star})$ and we have an edge $(e_{i}^{\star},e_{j}^{\star})$ in $\mathbf{L}(H)$ if the corresponding hyperedges intersect in $H$, i.e. $e_{i}\cap e_{j}\neq\emptyset$. Similarly to the clique expansion, many hypergraphs share the same line expansion, see Figure 1. However, joint knowledge of both the line and clique expansions seems to capture most of the structural information of the hypergraph. Indeed in Figure 1, hypergraphs $H_{1}$, $H_{2}$ and $H_{3}$ all share the same clique expansion while having different line expansions (they even have different node sets). Similarly to the clique graph, the adjacency matrix $A_{L}$ of the line graph can be derived from the incidence matrix via $A_{L}=\mathbf{I}^{T}\mathbf{I}-D_{e}$ where $D_{e}$ is the diagonal matrix of edge-lengths, with entries $\delta_{i}=|e_{i}|$. Let us define the dual hypergraph $H^{\star}=(E^{\star},V^{\star})$ as the hypergraph obtained by swapping the edge and node sets, i.e. $E^{\star}=(e_{1}^{\star},\ldots,e_{m}^{\star})$ and $V^{\star}=(v_{1}^{\star},\ldots v_{n}^{\star})$ where $v_{i}^{\star}=\{e_{j}^{\star}|v_{i}\in e_{j}\}$. We can easily observe, see Zhou et al., (2022), that the clique expansion is the line graph of the dual hypergraph: $\mathbf{C}(H)=\mathbf{L}(H^{\star})$.
@@
We can wonder if the joint knowledge of both the clique graph and the line graph allows to reconstruct the hypergraph. The answer is negative, as there exists pairs of distinct 0-1 matrices $(A,B)$ such that $AA^{T}=BB^{T}$ and $A^{T}A=B^{T}B$. Such matrices are called Gram mates, and we refer to Kim and Kirkland, (2022) for a proof of this result and a thorough analysis of families of matrices that generate Gram mates. However, it has been demonstrated in Corollary 1.1.1 of Kirkland, (2018) that if we uniformly pick two $m\times n$ 0-1 matrices, the probability that they are Gram mates decays exponentially as $(m,n)\to\infty$. Thus, for practical applications, the information loss from replacing a hypergraph with its clique and line expansions together is negligible. This representation also simplifies hypergraph analysis significantly.
@@
We consider hypergraphs where nodes are associated with discrete labels $\{1,\ldots,K\}$, representing communities or clusters. Intuitively, such hypergraphs exhibit a community structure, with denser connections between nodes sharing the same label compared to those with different labels. We assume a fully unsupervised setting, where we try to infer the labels (up to permutation) simply via the hypergraph structure. This generalization of the graph partitioning problem has been studied extensively, and we refer to Çatalyürek et al., (2023) for a survey of different approaches to tackle this problem. A popular method to determine the quality of a partitioning is given by modularity maximization, initially defined by Newman, (2006). We recall its definition in the graph case. For a graph $G=(V,E)$ and a partitioning $\mathbf{\mathcal{C}}=(C_{1},\ldots,C_{k})$ of $V$ into $k$ subsets, we define the modularity function $Q$ as:
$Q(\mathbf{\mathcal{C}})=\sum_{i=1}^{k}\frac{e_{G}(C_{i})}{|E|}-\sum_{i=1}^{k}\left(\frac{vol(C_{i})}{vol(V)}\right)^{2},$
where $e_{G}(C_{i})$ is the number of edges in the subgraph of $G$ induced by a node set $C_{i}$, and $vol(A)=\sum_{v\in A}deg(v)$ is the volume of any subset $A$ of nodes. Up to renormalization, the modularity function computes the difference between the number of edges uncut by the partitioning and the expected number of edges the same partitioning would yield in the Chung-Lu random graph model introduced in Chung and Lu, (2006). Maximizing $Q$ over every partitioning yields partitions that minimize cuts within communities, aligning with the underlying structure. There have been several works generalizing modularity for hypergraphs, notably Kumar et al., (2020) and Kamiński et al., (2024), by generalizing the Chung-Lu random model to hypergraphs, and allowing different weights to hyperedges of different sizes. Modularity also serves as a ground truth-free measure of the quality of a given clustering.
@@ 2.2 Ricci curvature and flows on graphs
This paper addresses the clustering problem using an extension of Ollivier-Ricci curvature to hypergraphs. This method extends prior works on Ollivier-Ricci curvature, initially defined in Ollivier, (2007) on metric spaces, and extended to the case of graphs in Lin et al., (2011). In Ni et al., (2019), the Ricci curvature is turned into a discrete flow such that edges between communities are heavily curved. They directly derive a clustering algorithm. We start by recalling the key concepts of their method in the case of graphs.
@@
Let $G=(V,E)$ be a weighted graph with a weight function $w:E\to\mathbb{R}_{\geq 0}$. In what follows, the function $w$ is set to be a dissimilarity measure between nodes. Let $d$ denote the shortest path distance between nodes induced by weights $w$. We adopt the framework introduced in Ni et al., (2019): for each node $x$, we define a probability measure $\nu_{x}^{\alpha,p}$ over $x$ and its neighbors given by:
$\nu_{x}^{\alpha,p}(y)=\begin{cases}\alpha&\text{if }y=x,\\ \frac{1-\alpha}{C}\exp(-d(x,y)^{p})&\text{if }y\in\mathcal{N}(x),\\ 0&\text{ otherwise, }\end{cases}$ (1)
where $\mathcal{N}(x)$ denotes the set of neighbors of $x$. Here, $C=\sum_{y\in\mathcal{N}(x)}\exp(-d(x,y)^{p})$ is a normalization constant to ensure that $\nu_{x}^{\alpha,p}$ is a probability measure. The parameters $\alpha\in[0,1]$ and $p$ respectively denote the amount of mass retained at $x$ and how much we want to penalize far neighbors. When not required or clear from context, we will drop the superscript and simply write $\nu_{x}^{\alpha,p}=\nu_{x}$.
@@
where $W$ is the 1-Wasserstein distance between probability measures for the cost function $d$, see Santambrogio, (2015). More precisely, a discrete transport plan $T$ between the measures $\nu_{x}$ and $\nu_{y}$ is an application $T:V\times V\to[0,1]$ such that $\sum_{v^{\prime}\in V}T(u,v^{\prime})=\nu_{x}(u)$ and $\sum_{u^{\prime}\in V}T(u^{\prime},v)=\nu_{y}(v).$ The 1-Wasserstein distance corresponds to the total cost of moving $\nu_{x}$ to $\nu_{y}$ with the optimal transport plan for a cost function $d$, i.e.
$W(\nu_{x},\nu_{y})=\inf\left\{\sum_{u,v\in V}T(u,v)d(u,v)|T\text{ transport plan between }\nu_{x}\text{ and }\nu_{y}\right\}.$
The sign of $\kappa$ provides information on the density of connections. Indeed, two nodes $x$ and $y$ belonging to the same community will typically share many neighbors, hence transporting $\nu_{x}$ to $\nu_{y}$ has a smaller cost than moving $x$ to $y$ and therefore induces a positive curvature. Conversely, nodes belonging to different communities do not share many neighbors, such that moving $\nu_{x}$ to $\nu_{y}$ requires traveling through the edge $(x,y)$ bridging the two communities. This in turn corresponds to a negative curvature. We refer to Figure 2 for a concrete example: when transporting $\nu_{v_{i}}$ to $\nu_{v_{j}}$ and $\nu_{u_{i}}$ to $\nu_{u_{j}}$, most of the mass remains on the nodes such that measures are transported at a low cost. In the extreme case where $\alpha=1/5$, all the mass remains and $\kappa(v_{i},v_{j})=\kappa(u_{i},u_{j})=1$. When transporting $\nu_{v_{1}}$ to $\nu_{u_{1}}$, all the $v_{i}$’s must be transported to $u_{i}$’s through the edge $(u_{1},v_{1})$ such that the total transport cost is larger than $d(u_{1},v_{1})$ and therefore $\kappa(u_{1},v_{1})<0$.
@@
Trimming edges of low curvature typically removes spurious connections between densely connected areas. Consequently, we can obtain a new graph where each connected component corresponds to a different community. We note that Ricci curvature can theoretically be computed for any pair of nodes in the graph. However, for our purposes, we only need to compute it for adjacent nodes.
@@
where $\kappa^{(l)}$ is the Ricci curvature computed with the distance induced by the weights $w^{(l)}$. The updated weights after the $l$-th iteration are referred to as the Ricci flow. This flow dynamic stretches edges of low curvature, making it more and more costly to transport measures between communities. The complete clustering algorithm is described by Algorithm 1, and simply trims edges with a high Ricci flow. The choice of the number of flow iterations $N$ is mostly heuristic, we refer to Ni et al., (2019) and the corresponding python library GraphRicciCurvature for some guidelines. The threshold $\tau$ at which to trim the weights is a crucial parameter and can also be taken as the one that maximizes the modularity function defined in Section 2.1, which evaluates clustering quality without ground truth labels.
Input: Graph $G=(V,E)$, Number of iterations NN, Threshold $\tau$
@@
This approach ensures that inter-community connections are down-weighted while preserving intra-community coherence. This enables a robust and interpretable method for clustering. Indeed, the method relies on direct computations of the Ollivier-Ricci curvature which is strongly associated to an intrinsic community partitioning of the graph, see Figure 2. We defer an extended comparison of the interpretability of Ricci-based clustering as opposed to neural network embeddings to Section 4.2.